# Gravity as a topological gauge theory

### v3

@@ 2.3 Associated action
Thus, by building an action from a linear combination of the Euler and Pontrjagin numbers associated to a manifold with Lorentzian Cartan geometry we retrieve the Holst action with a bare cosmological constant $\Lambda_{0}$ as well as the Euler, Pontrjagin and Nieh-Yan densities of the curvature $R$ and torsion $T$. According to [11], these three last terms should be present in any quantum theory of gravity.
@@
The main difference when comparing with the litterature is that in our case the coupling constants of the Euler and Pontrjagin densities are determined by the bare cosmological constant and by respectively the coupling constant of the Palatini action and the Barbero-Immirzi parameter as well as $r$. These dependencies on $G,\gamma$, $\Lambda_{0}$ and $r$ are interesting since a quantum version of the theory may yield effects sensitive to the bare cosmological constant $\Lambda_{0}$ in particular. We believe that probing the bare cosmological constant this way could be highly beneficial in constraining the vacuum contribution to the cosmological constant $\Lambda_{\text{vac}}$ since the two cases $\Lambda_{\text{vac}}=0$ and $\Lambda_{\text{vac}}=8\pi G\rho_{vac}$ bring about huge differences in the numerical value of $\Lambda_{0}$.
@@
One can also see that the coupling constant of the Euler density $-\dfrac{3\alpha_{1}}{4\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}$ obtained via the equation of motion (10) for the tetrad $\beta$ matches perfectly with what has been described in [12, section 3]. And in case $r=y=0$, it makes the action invariant under the duality symmetry consisting of the exchange $R^{ab}\leftrightarrow\dfrac{\Lambda_{0}}{3}\beta^{a}\wedge\beta^{b}$. The case of a dynamical $\ell$ will be studied in a forthcoming paper.
Also, in our case, properties of Characteristic classes imply that the choice of connection has no impact on the gauge action up to integrals of exact terms.
@@
We can consider $e=\dfrac{4\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell^{2}=-\dfrac{6k}{\Lambda_{0}}$ as fixed. Identifying the B-I parameter as $\gamma=\dfrac{2e}{3r+4y}$ then only leaves 2 degrees of freedom corresponding to the relative value of the parameters $r$ and $y$ and to $|k|$ the absolute value of $k$.
@@
The Pontrjagin density and other terms described here correspond to the Chern-Simons modified gravity [13] terms with a constant scalar field $\vartheta=(\dfrac{Gr}{\pi}-\dfrac{3}{\Lambda_{0}\gamma})\dfrac{\alpha_{1}}{2}$ and potential term for $\vartheta$, ${\displaystyle S_{\vartheta}=\int_{\mathscr{M}}V(\vartheta)}$, with :
$V(\vartheta)$ $=(\dfrac{1}{\gamma}-\dfrac{\Lambda_{0}Gr}{3\pi})\alpha_{1}R^{ab}\wedge\beta_{a}\wedge\beta_{b}\overbrace{-\dfrac{\Lambda_{0}\alpha_{1}}{12}\varepsilon_{abcd}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}}^{Cosmological~{}constant}$ (17)
$\qquad\overbrace{-\dfrac{3\alpha_{1}}{4\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{2\Lambda_{0}G\alpha_{1}r}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{i}\wedge\beta_{j}\bigr{)}}^{Nieh-Yan}.$
The Chern-Simons term also occurs in Loop Quantum Gravity (LQG) and String theory [13].
@@
$\left\{\begin{array}[]{ll}k=-1&\Leftrightarrow G=SO(4,1)\text{ for }\Lambda_{0}>0\\[5.69054pt] \left\{\begin{array}[]{ll}k=\pm 1\Leftrightarrow G=SO(4,1),SO(3,2)\\ \ell\rightarrow\infty\\ ¨r=y=0\end{array}\right\}\Leftrightarrow\left\{\begin{array}[]{ll}\Lambda_{0}\rightarrow 0\\ \gamma\rightarrow\infty\end{array}\right\}&\simeq\text{Einstein-Cartan}\\ k=1&\Leftrightarrow G=SO(3,2)\text{ for }\Lambda_{0}<0\end{array}\right.$ (26)
Taking $r\neq 0$ in the action (8) (with Pontrjagin number $\mathdutchcal{P}$ relative to the Cartan geometry $G/H$) secures the presence of the Barbero-Immirzi part of the action, together with the Pontrjagin density of $R$ and the Nieh-Yan term. On the other hand $y\neq 0$ is only linked to the Pontryagin and Barbero-Immirzi part of the Holst action.
3 Characteristic polynomials and action for the Möbius group $G=SO(4,2)/\{\pm I\}$
@@
and $H=CO(4,2)\ltimes\mathbb{R}^{*3,1}$ the maximal normal subgroup of $G$. A group element of H admits the matrix representation:
$H\ni h=\overbrace{\begin{pmatrix}\zeta&0&0\\ 0&\Lambda&0\\ 0&0&\zeta^{-1}\end{pmatrix}}^{\in G_{0}=CO(4,2)}\overbrace{\begin{pmatrix}1&\bar{u}&\tfrac{1}{2}\bar{u}u\\ 0&\mathds{1}&u\\ 0&0&1\end{pmatrix}}^{\in G_{1}=\mathbb{R}^{*3,1}}\text{ where }\Lambda\in SO(3,1),\Lambda^{\text{T}}\eta\Lambda=\eta\;;\;\zeta\Lambda\in CO(4,2)$
$u\in\mathbb{R}^{3,1}\;;\;\bar{u}=u^{\text{T}}\eta\in{\mathbb{R}^{3,1}}^{*}\text{ and }\zeta\in\mathbb{R}\setminus\{0\}.$
Its Lie algebra is $\mathfrak{g}=\mathfrak{so}(4,2)=\mathfrak{g}_{-1}\oplus\mathfrak{g}_{0}\oplus\mathfrak{g}_{1}$ with $\mathfrak{g}_{-1}\simeq\mathbb{R}^{3,1}$, $\mathfrak{g}_{0}\simeq\mathfrak{co}(4,2)=\mathfrak{so}(3,1)\oplus\mathbb{R}$ and $\mathfrak{g}_{1}\simeq\mathbb{R}^{*3,1}$. $\mathfrak{g}$ turns out to be a graded Lie algebra [2], namely,
@@
The generators of the above graded algebra are the same as the ones described previously in (2.1) except for $\mathfrak{g}_{0}\simeq\mathfrak{co}(4,2)$ which is generated by the $\{J_{ab}\}$’s and $\mathds{1}_{4}$. An element of $\mathfrak{g}_{0}\simeq\mathfrak{co}(4,2)$ is $\gamma_{\mathfrak{co}}=\frac{1}{2}\gamma_{\mathfrak{so}}^{ab}J_{ab}-z\mathds{1}_{4}$ with $\gamma_{\mathfrak{so}}\in\mathfrak{so}(3,,1)$ and $z\in\mathbb{R}$.
3.1 The Möbius algebra as a symmetric space
@@
The graded Lie algebra $\mathfrak{g}=\overbrace{\mathfrak{g}_{1}\oplus\mathfrak{g}_{0}}^{\mathfrak{h}}\oplus\overbrace{\mathfrak{g}_{-1}}^{\mathfrak{p}}\simeq{\mathbb{R}^{3,1}}^{*}\oplus\overbrace{\mathfrak{co}(4,2)}^{\mathfrak{so}(3,1)\oplus\mathbb{R}}\oplus\mathbb{R}^{3,1}$ of the Möbius group is reductive and is also a symmetric space $\mathfrak{g}=\mathfrak{co}(4,2)\oplus\mathfrak{m}$ where $\mathfrak{m}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{-1}\simeq{\mathbb{R}^{3,1}}^{*}\oplus\mathbb{R}^{3,1}$. The commutation relations (see (LABEL:eq:gradedLa) and [6, Example 5.2, Chapter XI]) are given by:
$[\mathfrak{g}_{0},\mathfrak{g}_{0}]\subset\mathfrak{g}_{0},\qquad[\mathfrak{g}_{0},\mathfrak{m}]\subset\mathfrak{m},\qquad[\mathfrak{m},\mathfrak{m}]\subset\mathfrak{g}_{0}\ .$ (1)
Let $\varphi=\varphi_{\mathfrak{g}_{1}}\oplus\varphi_{\mathfrak{g}_{0}}\oplus\varphi_{\mathfrak{g}_{-1}}=\bar{a}\oplus(c-z\mathds{1}_{4})\oplus b\in\mathfrak{g}$ with $a,b\in\mathbb{R}^{3,1}$, $c\in\mathfrak{so}(3,1)$, $z\in\mathbb{R}$ and $\bar{a}=\eta a^{T}$; $\bar{b}=\eta b^{T}\in\mathbb{R}^{*3,1}$. A matrix representation for $\mathfrak{g}$ is given by
$\varphi=\begin{pmatrix}z&\bar{a}&0\\ b&c&a\\ 0&\bar{b}&-z\end{pmatrix}=\begin{pmatrix}0&\bar{a}&0\\ 0&0&a\\ 0&0&0\end{pmatrix}+\begin{pmatrix}z&0&0\\ 0&c&0\\ 0&0&-z\end{pmatrix}+\begin{pmatrix}0&0&0\\ b&0&0\\ 0&\bar{b}&0\end{pmatrix}$ (2)
and the Lie bracket on $\mathfrak{g}$ is thus given by the commutator:
$[\varphi,\varphi^{\prime}]$ $=\begin{bmatrix}\begin{pmatrix}z&\bar{a}&0\\ b&c&a\\ 0&\bar{b}&-z\end{pmatrix},\begin{pmatrix}z^{\prime}&\bar{a}^{\prime}&0\\ b^{\prime}&c^{\prime}&a^{\prime}\\ 0&\bar{b}^{\prime}&-z^{\prime}\end{pmatrix}\end{bmatrix}$ (3)
$=\begin{pmatrix}\bar{a}b^{\prime}-\bar{a}^{\prime}b&z\bar{a}^{\prime}-z^{\prime}\bar{a}+\bar{a}c^{\prime}-\bar{a}^{\prime}c&\bar{a}a^{\prime}-\bar{a}^{\prime}a\\ z^{\prime}b-zb^{\prime}+cb^{\prime}-c^{\prime}b&[c,c^{\prime}]+b\bar{a}^{\prime}-b^{\prime}\bar{a}+a\bar{b}^{\prime}-a^{\prime}\bar{b}&ca^{\prime}-c^{\prime}a-z^{\prime}a+za^{\prime}\\ \bar{b}b^{\prime}-\bar{b}^{\prime}b&\bar{b}c^{\prime}-\bar{b}^{\prime}c-z\bar{b}^{\prime}+z^{\prime}\bar{b}&\bar{b}a^{\prime}-\bar{b}^{\prime}a\end{pmatrix}$
$=\begin{pmatrix}\bar{a}b^{\prime}-\bar{a}^{\prime}b&z\bar{a}^{\prime}-z^{\prime}\bar{a}+\bar{a}c^{\prime}-\bar{a}^{\prime}c&0\\ z^{\prime}b-zb^{\prime}+cb^{\prime}-c^{\prime}b&[c,c^{\prime}]+b\bar{a}^{\prime}-b^{\prime}\bar{a}+a\bar{b}^{\prime}-a^{\prime}\bar{b}&ca^{\prime}-c^{\prime}a-z^{\prime}a+za^{\prime}\\ 0&\bar{b}c^{\prime}-\bar{b}^{\prime}c-z\bar{b}^{\prime}+z^{\prime}\bar{b}&\bar{b}a^{\prime}-\bar{b}^{\prime}a\end{pmatrix}.$

### v5

@@ 2.3 Associated action
Thus, by building an action from a linear combination of these invariant polynomials and the Pontrjagin number associated to a manifold with Lorentzian Cartan geometry we retrieve the Holst action with a bare cosmological constant $\Lambda_{0}$ as well as the Euler, Pontrjagin and Nieh-Yan densities of the curvature $R$ and torsion $T$. According to [24], these three last terms should be present in any quantum theory of gravity.
The main difference compared to the literature is that in our case the coupling constants of the Euler and Pontrjagin densities are determined by the bare cosmological constant and by, respectively, the coupling constant of the Palatini action and the Barbero-Immirzi (B-I) parameter as well as $r$. Moreover, notice that one may identify the coefficient $\alpha_{3}=\dfrac{r+y}{8\pi^{2}}$ related to the Pontrjagin density built from $R$ with the parameter $\beta$ of [19, eq.(26)]. In this case one would then be able to identify (up to a sign convention) the last remaining physical parameter:
$r=-\dfrac{3\pi}{2\Lambda_{0}G}(\dfrac{1}{\gamma}\pm 1).$ (16)
@@
These dependencies on $G,\gamma$, $\Lambda_{0}$ and $r$ are interesting since a quantum version of the theory may yield effects sensitive to the bare cosmological constant $\Lambda_{0}$ in particular. We believe that probing the bare cosmological constant in this way could be highly beneficial in constraining the vacuum contribution to the cosmological constant $\Lambda_{\text{vac}}$ since the two cases $\Lambda_{\text{vac}}=0$ and $\Lambda=\Lambda_{\text{vac}}=8\pi G\rho_{vac}$ bring about huge differences in the numerical value of $\Lambda_{0}$.
@@
One can also see that the coupling constant of the Euler density $-\dfrac{3\alpha_{1}}{4\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}$ obtained via the equation of motion (9) for the tetrad $\beta$ matches perfectly with what has been described in [12, section III] such that the gauge action $S_{G}$ (15) is invariant under the "duality symmetry" consisting of the exchange $R^{ab}\leftrightarrow\dfrac{\Lambda_{0}}{3}\beta^{a}\wedge\beta^{b}$. It is worthwhile to notice that since by construction $S_{G}$ is quadratic in the curvature, $S_{G}$ turns out to be invariant under the (more general) duality symmetry $\bar{\Omega}\leftrightarrow-\bar{\Omega}$ which splits as:
$\begin{cases}\bar{\Omega}_{\mathfrak{h}}\leftrightarrow-\bar{\Omega}_{\mathfrak{h}}\Rightarrow R\leftrightarrow\dfrac{\Lambda_{0}}{3}\beta\wedge\bar{\beta}\\ \bar{\Omega}_{\mathfrak{m}}\leftrightarrow-\bar{\Omega}_{\mathfrak{m}}\Rightarrow T/\ell\leftrightarrow-T/\ell\ .\end{cases}$ (17)
and reduces to the case of [12] upon solely considering the $\mathfrak{h}$-sector of the curvature. Remark that $S_{G}$ is invariant under each of these transformations independently.
In addition, a direct comparison with the two perturbative expansion parameters $\alpha$,$\beta$ given in [18] yields:
$\alpha$ $=\dfrac{G\Lambda_{0}}{3(1-\gamma^{2})}=-\dfrac{r+2y}{e}\beta=-\dfrac{\pi}{2e}\dfrac{(r+2y)^{2}}{(r+2y)^{2}-e^{2}}$
$\beta$ $=\dfrac{\gamma G\Lambda_{0}}{3(1-\gamma^{2})}=\dfrac{\pi}{2}\dfrac{r+2y}{(r+2y)^{2}-e^{2}}.$
As for $k$, the case of a dynamical $\ell$ will be studied in a forthcoming paper.
@@
We can now consider $e=\dfrac{8\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell^{2}=-\dfrac{3k}{\Lambda_{0}}$ as fixed. Upon identifying the B-I parameter as $\gamma=\dfrac{e}{r+2y}$, only two degrees of freedom corresponding to the relative value of the parameters $r$ and $y$ and to $|k|$ (the absolute value of $k$) are left.
@@
The Pontrjagin density and other terms described in (15) correspond to the Chern-Simons modified gravity [25] terms with a constant scalar field $\vartheta=-\big{(}\dfrac{2Gr}{\pi}+\dfrac{3}{\Lambda_{0}\gamma}\big{)}\dfrac{\alpha_{1}}{2}$ and potential term for $\vartheta$, ${\displaystyle S_{\vartheta}=\int_{\mathscr{M}}V(\vartheta)}$, with:
$V(\vartheta)$ $=(\dfrac{1}{\gamma}-\dfrac{2\Lambda_{0}Gr}{3\pi})\alpha_{1}R^{ab}\wedge\beta_{a}\wedge\beta_{b}\overbrace{-\dfrac{\Lambda_{0}\alpha_{1}}{12}\varepsilon_{abcd}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}}^{Cosmological~{}constant}$ (18)
$\qquad\overbrace{-\dfrac{3\alpha_{1}}{4\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{4\Lambda_{0}G\alpha_{1}r}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}.$
The Chern-Simons term also occurs in Loop Quantum Gravity (LQG) and String theory as mentionned in [25].
@@
$\left\{\begin{array}[]{ll}k=-1&\Leftrightarrow G=SO(4,1)\text{ for }\Lambda_{0}>0\\[5.69054pt] \left\{\begin{array}[]{ll}k=\pm 1\Leftrightarrow G=SO(4,1),SO(3,2)\\ \ell\rightarrow\infty\\ r=-2y=0\end{array}\right\}\Leftrightarrow\left\{\begin{array}[]{ll}\Lambda_{0}\rightarrow 0\\ \gamma\rightarrow\infty\end{array}\right\}&\simeq\text{Einstein-Cartan}\\ k=1&\Leftrightarrow G=SO(3,2)\text{ for }\Lambda_{0}<0\end{array}\right.$ (27)
Taking $ry\neq 0$ and $r\neq-2y$ in the action (8) secures the presence of the Barbero-Immirzi part of the Holst action, together with the Pontrjagin density of $R$, the Nieh-Yan term and a non diverging Barbero-Immirzi parameter.
3 Invariant polynomials and action for the Möbius group $G=SO(4,2)/\{\pm I\}$
@@
and $K=\overbrace{CO(3,1)}^{H}\ltimes\mathbb{R}^{3,1*}$ the maximal normal subgroup of $G$. A group element of K admits the matrix representation:
$K\ni\overbrace{\begin{pmatrix}\zeta&0&0\\ 0&\Lambda&0\\ 0&0&\zeta^{-1}\end{pmatrix}}^{\in H=CO(3,1)}\overbrace{\begin{pmatrix}1&\bar{u}&\tfrac{1}{2}\bar{u}u\\ 0&\mathds{1}&u\\ 0&0&1\end{pmatrix}}^{\in\mathbb{R}^{3,1*}}\text{ where }\begin{array}[]{ll}\Lambda\in SO(3,1),\Lambda^{\text{T}}\eta\Lambda=\eta\;;\;\zeta\Lambda\in CO(3,1),\\ u\in\mathbb{R}^{3,1}\;;\;\bar{u}=u^{\text{T}}\eta\in{\mathbb{R}^{3,1}}^{*}\text{ and }\zeta\in\mathbb{R}^{+}\end{array}$
where $\mathbb{R}^{+}$ denotes the multiplicative group of positive real numbers.
The Lie algebra of $G$ is $\mathfrak{g}=\mathfrak{so}(4,2)=\mathfrak{g}_{-1}\oplus\mathfrak{g}_{0}\oplus\mathfrak{g}_{1}$ with $\mathfrak{g}_{-1}\simeq\mathbb{R}^{3,1}$, $\mathfrak{g}_{0}\simeq\mathfrak{co}(3,1)=\mathfrak{so}(3,1)\oplus\mathbb{R}$ and $\mathfrak{g}_{1}\simeq\mathbb{R}^{3,1*}$. The Lie algebra $\mathfrak{g}$ turns out to be a graded Lie algebra [2], namely,
@@
The generators of the above graded algebra are the same as the ones described previously in (2.1) except for $\mathfrak{g}_{0}\simeq\mathfrak{co}(3,1)$ which is generated by the $\{J_{ab}\}$’s and $\mathds{1}_{4}$. An element of $\mathfrak{co}(3,1)\simeq\mathfrak{g}_{0}$ is $\gamma_{\mathfrak{co}}=\frac{1}{2}\gamma_{\mathfrak{so}}^{ab}J_{ab}-z\mathds{1}_{4}$ with $\gamma_{\mathfrak{so}}\in\mathfrak{so}(3,,1)$ and $z\in\mathbb{R}$.
3.1 The Möbius Lie algebra as a symmetric Lie algebra
Let us consider $H=CO(3,1)$ as structure group for the Cartan geometry to be discussed in this part. This implies a geometry that differs from the whole isotropic Möbius group as introduced in [11, Lemma 1.8,p.269]). It rather corresponds to $SO(3,1)\times\mathbb{R}^{+}$ (Lorentz×$\times$×Weyl) the isotropic part of the Weyl model which can be embedded into the Möbius model, see [11, middle p.277].
The following decomposition of the graded Lie algebra
$\mathfrak{g}=\overbrace{\mathfrak{g}_{0}}^{\mathfrak{h}}\oplus\overbrace{\mathfrak{g}_{1}\oplus\mathfrak{g}_{-1}}^{\mathfrak{m}}\simeq\overbrace{\mathfrak{co}(3,1)}^{\mathfrak{so}(3,1)\oplus\mathbb{R}}\oplus{\mathbb{R}^{3,1}}^{*}\oplus\mathbb{R}^{3,1}$ (1)
@@ 3.1 The Möbius Lie algebra as a symmetric Lie algebra
of the Möbius group is reductive with respect to $H$ and is also a symmetric Lie algebra $\mathfrak{g}=\mathfrak{co}(3,1)\oplus\mathfrak{m}$ where $\mathfrak{m}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{-1}\simeq{\mathbb{R}^{3,1}}^{*}\oplus\mathbb{R}^{3,1}$. The commutation relations (see (LABEL:eq:gradedLa) and [14, Example 5.2 and Chapter XI]) are given by
$[\mathfrak{g}_{0},\mathfrak{g}_{0}]\subset\mathfrak{g}_{0},\qquad[\mathfrak{g}_{0},\mathfrak{m}]\subset\mathfrak{m},\qquad[\mathfrak{m},\mathfrak{m}]\subset\mathfrak{g}_{0}$ (2)
making $\mathfrak{g}=\mathfrak{g}_{0}\oplus\mathfrak{m}$ a symmetric Lie algebra.
According to the graded Lie algebra decomposition of $\mathfrak{g}$, let
$\varphi_{\mathfrak{g}}=\varphi_{1}\oplus\varphi_{0}\oplus\varphi_{-1}\in\mathfrak{g}=\mathfrak{g}_{1}\,\oplus\,\mathfrak{g}_{0}\,\oplus\,\mathfrak{g}_{-1}$
for which we adopt the matrix presentation [^14] Accordingly, the Killing form on $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ is computed to be $K_{\mathfrak{g}}(\varphi,\varphi^{\prime})=\operatorname{Tr}(ad_{\varphi}\circ ad_{\varphi^{\prime}})=n\operatorname{Tr}(cc^{\prime})+2n\big{(}zz^{\prime}+\eta(b,a^{\prime})+\eta(a,b^{\prime})\big{)}$ where $n$ corresponds to the one coming from $\dim\mathfrak{so}(n-1,1)=n(n-1)/2$.
$\varphi_{\mathfrak{g}}=\varphi_{1}\oplus\varphi_{0}\oplus\varphi_{-1}=\begin{pmatrix}z&\bar{a}&0\\ b&c&a\\ 0&\bar{b}&-z\end{pmatrix}=\begin{pmatrix}0&\bar{a}&0\\ 0&0&a\\ 0&0&0\end{pmatrix}+\begin{pmatrix}z&0&0\\ 0&c&0\\ 0&0&-z\end{pmatrix}+\begin{pmatrix}0&0&0\\ b&0&0\\ 0&\bar{b}&0\end{pmatrix}$ (3)
with $a,b\in\mathbb{R}^{3,1}$, $c\in\mathfrak{so}(3,1)$, $z\in\mathbb{R}$ and $\bar{a}=a^{T}\eta,\bar{b}=b^{T}\eta\in\mathbb{R}^{*3,1}$. It is worthwhile to notice that $E=\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&-1\end{pmatrix}\in\mathfrak{g}_{0}$ the generator for dilation defines the grading of $\mathfrak{g}$ by $\mathfrak{g}_{j}=\{\varphi\in\mathfrak{g}\text{ s.t. }[E,\varphi]=j\varphi\},j=0,\pm 1$. Moreover, through the isomorphism $\mathfrak{g}\simeq{\mathbb{R}^{3,1}}^{*}\oplus\mathfrak{co}(3,1)\oplus\mathbb{R}^{3,1}$, one can also write [2]
$\varphi_{\mathfrak{g}}=\varphi_{1}\oplus\varphi_{\mathfrak{co}}\oplus\varphi_{-1}=\varphi_{1}\oplus(\varphi_{\mathfrak{so}}-\varphi_{\mathbb{R}}\mathds{1}_{4})\oplus\varphi_{-1}=\bar{a}\oplus(c-z\mathds{1}_{4})\oplus b.$
Let us now consider the mutation [^15] See Definition 2.1 given above. $\mathfrak{g}^{\prime}=\mu(\mathfrak{g})=\mu(\mathfrak{h}\oplus\mathfrak{m})=\mathfrak{h}\oplus\mu(\mathfrak{m})=\mathfrak{h}\oplus\mathfrak{m}^{\prime}$ given by:
$\varphi_{\mathfrak{g}}=\begin{pmatrix}z&\bar{a}&0\\ b&c&a\\ 0&\bar{b}&-z\end{pmatrix}\mapsto\varphi_{\mathfrak{g}^{\prime}}=\mu(\varphi_{\mathfrak{g}})=\begin{pmatrix}z&\bar{\gamma}_{2}\bar{a}&0\\ \gamma_{1}b&c&\gamma_{2}a\\ 0&\bar{\gamma}_{1}\bar{b}&-z\end{pmatrix}\in\mathfrak{g}^{\prime}$ (4)
where the scalar mutation parameters $\gamma_{1},\bar{\gamma}_{1}\gamma_{2},\bar{\gamma}_{2}$ are arranged such that $\mu$ is a linear isomorphism, in particular, $\gamma_{1}\bar{\gamma}_{2}=\bar{\gamma}_{1}\gamma_{2}$. Thus, it is merely a change of scale within the symmetric Lie algebra which preserves the splitting $\mathfrak{m}=\mathfrak{g}_{-1}\oplus\mathfrak{g}_{1}$ according to the graded structure.
Let us choose $\gamma_{1}=\bar{\gamma}_{1}=\dfrac{1}{\ell}$, $\gamma_{2}=\bar{\gamma}_{2}=\dfrac{1}{\ell^{\prime}}$ (where $\ell$ and $\ell^{\prime}$ are non-null real numbers) [^16] In that case, the Killing form on $\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}^{\prime}$ can be easily computed by performing the scale transformations $b\mapsto\gamma_{1}b$ and $a\mapsto\gamma_{2}a$ into the expression for $K_{\mathfrak{g}}$, so that $K_{\mathfrak{g}^{\prime}}(\varphi,\varphi^{\prime})=n\operatorname{Tr}(cc^{\prime})+2n\Bigl{(}zz^{\prime}+\gamma_{1}\gamma_{2}\bigl{(}\eta(b,a^{\prime})+\eta(a,b^{\prime})\bigr{)}\Bigr{)}.$
The Lie bracket on $\mathfrak{g}^{\prime}$ is thus given by the commutator:
$[\varphi,\varphi^{\prime}]$ $=\begin{bmatrix}\begin{pmatrix}z&\dfrac{1}{\ell^{\prime}}\bar{a}&0\\ \dfrac{1}{\ell}b&c&\dfrac{1}{\ell^{\prime}}a\\ 0&\dfrac{1}{\ell}\bar{b}&-z\end{pmatrix},\begin{pmatrix}z^{\prime}&\dfrac{1}{\ell^{\prime}}\bar{a}^{\prime}&0\\ \dfrac{1}{\ell}b^{\prime}&c^{\prime}&\dfrac{1}{\ell^{\prime}}a^{\prime}\\ 0&\dfrac{1}{\ell}\bar{b}^{\prime}&-z^{\prime}\end{pmatrix}\end{bmatrix}$ (5)
$=\begin{pmatrix}\dfrac{1}{\ell\ell^{\prime}}(\bar{a}b^{\prime}-\bar{a}^{\prime}b)&\dfrac{1}{\ell^{\prime}}(z\bar{a}^{\prime}-z^{\prime}\bar{a}+\bar{a}c^{\prime}-\bar{a}^{\prime}c)&\dfrac{1}{\ell^{\prime 2}}(\bar{a}a^{\prime}-\bar{a}^{\prime}a)\\ \dfrac{1}{\ell}(z^{\prime}b-zb^{\prime}+cb^{\prime}-c^{\prime}b)&[c,c^{\prime}]+\dfrac{1}{\ell\ell^{\prime}}(b\bar{a}^{\prime}-b^{\prime}\bar{a}+a\bar{b}^{\prime}-a^{\prime}\bar{b})&\dfrac{1}{\ell^{\prime}}(ca^{\prime}-c^{\prime}a-z^{\prime}a+za^{\prime})\\ \dfrac{1}{\ell^{2}}(\bar{b}b^{\prime}-\bar{b}^{\prime}b)&\dfrac{1}{\ell}(\bar{b}c^{\prime}-\bar{b}^{\prime}c-z\bar{b}^{\prime}+z^{\prime}\bar{b})&\dfrac{1}{\ell\ell^{\prime}}(\bar{b}a^{\prime}-\bar{b}^{\prime}a)\end{pmatrix}$
$=\begin{pmatrix}\dfrac{1}{\ell\ell^{\prime}}(\bar{a}b^{\prime}-\bar{a}^{\prime}b)&\dfrac{1}{\ell^{\prime}}(z\bar{a}^{\prime}-z^{\prime}\bar{a}+\bar{a}c^{\prime}-\bar{a}^{\prime}c)&0\\ \dfrac{1}{\ell}(z^{\prime}b-zb^{\prime}+cb^{\prime}-c^{\prime}b)&[c,c^{\prime}]+\dfrac{1}{\ell\ell^{\prime}}(b\bar{a}^{\prime}-b^{\prime}\bar{a}+a\bar{b}^{\prime}-a^{\prime}\bar{b})&\dfrac{1}{\ell^{\prime}}(ca^{\prime}-c^{\prime}a-z^{\prime}a+za^{\prime})\\ 0&\dfrac{1}{\ell}(\bar{b}c^{\prime}-\bar{b}^{\prime}c-z\bar{b}^{\prime}+z^{\prime}\bar{b})&\dfrac{1}{\ell\ell^{\prime}}(\bar{b}a^{\prime}-\bar{b}^{\prime}a)\end{pmatrix}$